The molar solubility of Mg(OH)2 3.72x10-5 M.  Will Mg(OH)2 precipitate when 100mL of 0.001M solution of MgCl2 mixed with 200mL of 0.002M KOH? (12pts)
Answer #1

Auswers mg cl2 199) + 2 Kom (aa) mg/04/215) + Excl(aa) mg109/2 15) might (aq) + 2041aa) Kap= 3.72410-Sm Indles of mach = vole